OpenStudy (anonymous):

determine if the following function is even p(x)=2x^4-6x+1

OpenStudy (anonymous):

hint, one of the exponents is NOT even

OpenStudy (anonymous):

your teacher wants you to write \[p(-x)=(-x)^2-6\times (-x)+1=x^2+6x+1\] and since this is NOT \[-p(x)\] the answer is no

OpenStudy (campbell_st):

for a function to be even the condition is f(-x) = f(x). for a function to be odd the condition is f(-x) = -f(x) if neither condition is satisfied, the function is neither odd nor even